Exercise 2-44 (LO2-2) A total of 4.8 percent of the prime time viewing audience watched shows on ABC, 6.7 percent watched shows on C Fox, 5.8 percent on NBC, 1.7 percent on Warner Brothers, and 1.9 percent on UPN. A total of shows on other cable networks, such as CNN and ESPN BS, 5.2 percent on What will be the prime time viewing percentage of the pie chart for the following data? (Round your answers to 2 decimal places) Prime time viewing ABC CBS Fox NBC Warner UPN
Answer #1

Answer:

Total prime time viewing for all channels =4.8+6.7+5.2+5.8+1.7+1.9=26.1

Prime time viewing percentage for ABC= prime time viewing for abc/ Total prime time viewing for channels=4.8/26.1=18.39%

So for other channels

Channel prime time viewing percentage
ABC 4.8/26.1=18.39%
CBS 6.7/26.1=25.67%
FOX 5.2/26.1=19.92%
NBC 5.8/26.1=22.22%
WARNER 1.7/26.1=6.51%
UPN

1.9/26.1=7.28%